# Word Definitions from Large Language Models

### v2

@@ 5.2 Word Frequency Analysis
To resolve this debate, we compare the average cosine distance between BERT-encoded dictionary and generated definitions, partitioned by class into words of high frequency, moderate frequency, and rare words. The results in Table 8 show words of middle frequency produce the best dictionary-generated for all models, across all dictionaries. The differences between frequency tiers is generally quite modest, generally less than the impact of model version. GPT-3.5 appears to generate slightly better definitions than the later GPT-4 by this metric, while the choice of prompt has little effect on the accurracy of resdult from either model.
@@
Table 8: Cosine distance based on BERT for different word frequency.
5.3 Are GPT-Generated Definitions Accurate?
Researchers have found that ChatGPT can easily generate answers with complete assurance, even sometimes the answer is wrong Lehnert (2023). To evaluate the correctness of GPT-generated definitions, we manually compare the definitions of the 50 words with the largest Euclidean distance between embedding definitions between Merriam-Webster and GPT-4.
We tabulate the words with the largest such distances in Table 9. Of the fifty words we manually evaluate, the biggest distances occur when models admit they do not know the definitions of generally obscure words (often proper names and abbreviations) that appeared in Merriam-Webster. Only two of the fifty words (cordier and imon) were unknown to GPT-4, compared to 16 words GPT-3 did not understand. Of the 48 words for which GPT-4 ventured a definition, we deemed all of them similar except for “Acton”, which denoted a place in GPT-4 and a person in Merriam-Webster. We conclude that the model generated definitions are generally of high quality, consistent with those from published dictionaries.
Word
Definition from GPT3
Definition from GPT4
Definition in Merriam-Webster
paget I’m sorry, as an AI language model, I cannot define the word "paget" accurately without any context. "Paget" is primarily known as a surname of Norman origin and can be seen in various contexts. Sir James 1814–1899 English surgeon and pathologist.
acton I’m sorry, but "acton" is not a word in the English language. "Acton" is primarily known as a place name, particularly in England. It’s a district in the West London, in the Borough of Ealing. 1st Baron 1834–1902 John Emerich Edward Dalberg-Acton, English historian.
nov I’m sorry, I cannot define the word "nov" as it does not exist or have any clear meaning in the English language. "Nov" is most commonly recognized as an abbreviation for the month of November. November.
cordier As an AI language model, I cannot determine the context of the word "cordier". I’m sorry, but as of my knowledge cutoff in September 2021, "cordier" doesn’t have a specific definition in English. It seems like it could be a surname. comparative of cordy.
inion There is no word "inion" in the English language. It may be a misspelling of "onion," which is a type of vegetable. The inion is a term used in anatomy that refers to the most prominent projection of the occipital bone at the lower rear part of the skull in humans and other hominids. the external occipital protuberance of the skull.
Table 9: Representative examples of the words whose GPT definition embeddings are far from online dictionary definition embedding.
6 Word Embeddings and Definitions
Word embeddings are vector representations that capture the semantics of word usage. What dictionary definitions are for people, word embeddings are for NLP models: an easily-understood representation of the meaning of a vocabulary word. In this section we will directly compare traditional word embeddings to text embeddings of explicit dictionary embeddings, to help establish the level of correspondence between these representations.
Specifically, we obtain GloVE and 300-dimensional FastText word embeddings on all the words, and compare the closest neighbors of a word based on definition and word embeddings. We cannot compare these embeddings directly, because the underlying spaces and even dimensionality are incomparable.
@@ 6 Word Embeddings and Definitions
6.1 Frequency Analysis
Table 10 compares published and generated definitions with popular word embeddings, GloVE and FastText, using the correlation measure defined above. We distinguish between our three tranches of words, high-frequency, moderate, and rare.
The results are consistent across all sources and embeddings: high-frequency words show better consistency between definition-word embeddings than moderate-frequency words, which are substantially better than low frequency words. As our results in Section 5.2 show the definitions of roughly equal quality across frequency tranches, the difference must be due to the word embeddings themselves becoming less accurate with diminishing usage frequency. This presumably results from embeddings that are trained on insufficient amounts of data.
@@ 6.1 Frequency Analysis
Table 10: Pearson correlation between word embedding and definition embedding based on cosine distance for different word frequency.
7 Conclusion and Future Work
The relationship between full text descriptions and embeddings is an interesting one. Our work demonstrates that model-generated definitions are accurate and lead to other useful representations.
Our experiments to detect long common phrases in alternate definitions of a particular word motivates questions of how long such phrases must be in generated text. How just how far apart can two texts be in surface form while still meaning the same thing? A study of situations where LLMs are highly constrained to specific word/phrase choices may send interesting light on the situation.
Limitations
The generative models powering ChatGPT are undergoing rapid evolution, so the observations in this paper cannot be guaranteed to hold in future releases. However, we do anticipate that the models will only continue to improve, and since our results show that they already exhibit similar accuracy and originality as published dictionaries we would be surprised of our observations do not continue to hold with future releases.
Having established basic agreement between publishing and generated definitions, as we have done in this paper, is a less demanding task than validating the completeness of multiple senses or alternate usages. We also have restricted our attention to English words and definitions. It would be interesting to evaluate model fidelity in other languages, or translation (e.g. “Define the French word rouge.”).
Appendix A Appendix
Model Definition of "wasp"
GPT3-1 a winged insect with a slender, often brightly colored body, the females of which have a poisonous sting
GPT3-2 A wasp is a predatory flying insect with a narrow waist and a sting, belonging to one of several families of the order Hymenoptera, such as the Vespidae, which includes the social wasps and hornets, or the Ichneumonidae, which includes parasitic wasps.
GPT4-1 "Wasp" has a couple of different meanings, depending on the context: 1. In entomology (the study of insects), a wasp is a type of flying insect, known for its black and yellow body, narrow waist, and ability to sting. There are many thousands of species of wasps, and they are found all over the world. 2. In social contexts, especially in the United States, WASP is an acronym for White Anglo-Saxon Protestant. It refers to a socio-economic and cultural group that has traditionally been dominant in American society, particularly among the upper class. This usage can sometimes carry implications about privilege and exclusivity. Remember that the context usually makes clear which meaning is intended.
GPT4-2 A wasp is a type of insect that is part of the order Hymenoptera and the suborder Apocrita, which is not a true ant or a bee. There are many different species of wasps, which can vary significantly in terms of size, color, and behavior. Most wasps are known for their ability to sting in self-defense. Unlike bees, which typically can sting only once before dying, many wasps can sting multiple times. Wasp stings are painful and can be dangerous to people who are allergic to them. Wasps can be categorized as either social or solitary. Social wasps, like the common yellowjacket, live in large colonies, and their nests can be found in trees, roof eaves, or underground. Solitary wasps, on the other hand, live alone and do not form colonies. Wasps play an important role in the ecosystem, as they prey on other pests, helping to control their populations. Some species are also important pollinators.
WordNet white Anglo-Saxon Protestant (a white person of Anglo-Saxon ancestry who belongs to a Protestant denomination); (social or solitary hymenopterans typically having a slender body with the abdomen attached by a narrow stalk and having a formidable sting)
Merriam-Webster any of numerous social or solitary winged hymenopterous insects…; any of various hymenopterous insects…; an American of Northern European and especially British ancestry and of Protestant background, especially : a member of the dominant and the most privileged class of people in the U.S.
Dictionary.com any of numerous social or solitary hymenopterous insects…; a person who is snappish or petulant; a member of the Women’s Air Force Service Pilots, an organization disbanded in 1944; a white Anglo-Saxon Protestant; a member of the privileged, established white upper middle class in the U.S.
Table 11: Definitions of the word “wasp” from GPT models and online dictionaries.

### v3

@@ V-B Word Frequency Analysis
To resolve this debate, we compare the average cosine distance between SBERT-encoded dictionary and generated definitions, partitioned by class into words of high frequency, moderate frequency, and rare words. The results in Table III show words of middle frequency produce the best dictionary-generated for all models, across all dictionaries. The differences between frequency tiers is generally quite modest, less than the impact of model version. GPT-3.5 appears to generate slightly better definitions than the later GPT-4 by this metric, while the choice of prompt has little effect on the accuracy of result from either model.
@@
TABLE III: Cosine distance based on SBERT for different word frequency.
V-C Are GPT-Generated Definitions Accurate?
Researchers have found that ChatGPT can easily generate answers with complete assurance, even sometimes the answer is wrong [25]. To evaluate the correctness of GPT-generated definitions, we manually compare the definitions of the 50 words with the largest Euclidean distance between embedding definitions between Merriam-Webster and GPT-4. Of these fifty words we manually evaluate, the biggest distances occur when models admit they do not know the definitions of generally obscure words (often proper names and abbreviations) that appeared in Merriam-Webster. Only two of the fifty words (“cordier” and “imon”) were unknown to GPT-4, compared to 16 words GPT-3 did not understand. Of the 48 words for which GPT-4 ventured a definition, we deemed all of them similar except for “Acton”, which denoted a place in GPT-4 and a person in Merriam-Webster. We conclude that the model generated definitions are generally of high quality, consistent with those from published dictionaries.
VI Word Embeddings and Definitions
Word embeddings are vector representations that capture the semantics of word usage. What dictionary definitions are for people, word embeddings are for NLP models: an easily-understood representation of the meaning of a vocabulary word. In this section we will directly compare traditional word embeddings to text embeddings of explicit dictionary embeddings, to help establish the level of correspondence between these representations. Specifically, we obtain GloVE and 300-dimensional FastText word embeddings on all the words, and compare the closest neighbors of a word based on definition and word embeddings. We cannot compare these embeddings directly, because the underlying spaces and even dimensionality are incomparable.
@@ VI Word Embeddings and Definitions
Table IV compares published and generated definitions with popular word embeddings, GloVE and FastText, using the correlation measure defined above. We distinguish between our three tranches of words, high-frequency, moderate, and rare.
The results are consistent across all sources and embeddings: high-frequency words show better consistency between definition-word embeddings than moderate-frequency words, which are substantially better than low frequency words. As our results in Section V-B show the definitions of roughly equal quality across frequency tranches, the difference must be due to the word embeddings themselves becoming less accurate with diminishing usage frequency. This presumably results from embeddings that are trained on insufficient amounts of data.
@@
TABLE IV: Pearson correlation between word embedding and definition embedding based on cosine distance for different word frequency.
VII Conclusion and Future Work
Our work demonstrates that while LLM-generated definitions differ from classical dictionary definitions in phrasing, they remain consistent and accurate in meaning. We also reveal that LLM-generated definitions are robust, particularly for low-frequency words, where they outperform traditional word embeddings. The scope of future work can be expanded to include multilingual definitions and domain-specific vocabularies to gain more insights into the potential and limitations of LLMs for semantic representation and language understanding.